OpenStudy (anonymous):

If the maximum value in the range of the function y=f(x) is 6, what is the maximum value in the range of the function y=3f(x-1)?

OpenStudy (anonymous):

either the question is 100% clear or u said the maximum is 6 , and asked what's the maximum? :\

OpenStudy (anonymous):

i know. i left off part of the question sorry

OpenStudy (anonymous):

it's okay @Megan2015 , now it's clear ;)

OpenStudy (anonymous):

Do you know how to do it? @Ahmad1

OpenStudy (anonymous):

for the function f(x-1) let y =x-1 , thus the maximum of function in term of is f(y)=6 since f(x) max=6 , so the maximum of 3*f(y)=3* maxium f(y) =3*6=18

OpenStudy (anonymous):

did that make sense ?

OpenStudy (anonymous):

not really so the maximum of the function is the maximum of the other function times 3 because there is a coefficient in the second problem?

OpenStudy (anonymous):

@Ahmad1

OpenStudy (anonymous):

yes , and the trick x-1 instead of x does not effect (if the domain is not restricted) imagine that the maximum occur at 0 such that f(0)=6 so the maximum in f(x-1) will occur at 1 such that f(1-1)=6 , then you multiply by 3

OpenStudy (anonymous):

oh ok! thank you!

OpenStudy (anonymous):

Another way of putting it is that \(f(x-1)\) is \(f(x)\) shifted horizontally, so the \(x-1\) shouldn't affect the maximum value.

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

@SithsAndGiggles by shifted do you mean to the left?

OpenStudy (anonymous):

For this particular problem the direction of the shift doesn't matter. But in general, given \(f(x-c)\), this means you have \(f(x)\) shifted to the right by \(c\) units if \(c>0\), or shifted to the left by \(c\) units if \(c<0\).

OpenStudy (anonymous):

@SithsAndGiggles okay. that makes sense. but the maximum value still changes right?

OpenStudy (anonymous):

No. Here's an example: